OpenStudy (anonymous):

Solve the equation 2(x + 4) = 30

OpenStudy (anonymous):

1. distribute 2x+8 = 30

OpenStudy (anonymous):

2x+8-8=30-8

OpenStudy (anonymous):

2x = 22

OpenStudy (anonymous):

x = 11